# Exploring Citation Diversity in Scholarly Literature: An Entropy-Based Approach

### v1

@@ Abstract
This study explores global citation diversity, examining its various patterns across countries and academic disciplines. We analyzed citation distributions in top institutes worldwide, revealing that the higher citation end of the distribution follow Power law or Pareto law pattern and the Pareto law’s scaling exponent changes with the number of institutes considered. An entropy based novel citation inequality measure has been introduced, enhancing the precision of our analysis. Our findings show that countries with small and large economies often group similarly based on citation diversity, with shifting the groupings as the number of institutes considered changes. Moreover, we analyzed citation diversity among award-winning scientists across six scientific disciplines, finding significant variations. We also explored the evolution of citation diversity over the past century across multiple fields. A gender-based study in various disciplines highlights citation inequalities among male and female scientists. Our innovative citation diversity measure stands out as a vital tool for evaluating citation inequality, providing insights beyond what traditional citation counts can offer. This thorough analysis deepens our understanding of global scientific contributions and promotes a more equitable view of academic accomplishments.
Keywords: Citation inequality; Diversity measure; General class of Entropy; Scholarly literature; Award winners.
@@ I Introduction
Citations are the currency of academia, reflecting impacts and influences of research publications. Ideally, a fair citation landscape would see recognition distributed proportionally to the quality and contribution of research. However, a growing concern in recent years is the phenomenon of citation inequality. This refers to the uneven distribution of citations, where a small number of papers garner a disproportionate share of citations but the vast majority receive significantly fewer [1, 2, 3]. Understanding the dynamics of citation inequality is crucial for several reasons. Firstly, it raises concerns about the fairness and effectiveness of the current evaluation system in academia. Overemphasis on citations can lead to a ‘publish or perish’ mentality, potentially favoring sensational or trendy research over groundbreaking discoveries that take longer to gain recognition. Secondly, citation inequality can hinder the dissemination and application of valuable research, particularly from less established scholars or emerging fields.
Measuring citation inequality is a challenging task that has garnered significant attention in the research community. Interestingly, different approaches have been employed to quantify these measurements, drawing parallels to methods used in assessing economic inequality [2, 4, 5]. The Hirsch index, or h-index, is also widely recognized as a tool for measuring citation inequality, as highlighted by multiple studies [6, 7, 8]. However, in the studies of citation inequality, the generalised entropy index, used in the area of income inequality [9], has not yet been adopted as a useful measure.
Motivated from all these works, here, for the first time, an entropy based diversity measure is used to study the citation inequality. This research article delves into the complexities of citation inequality through an innovative approach utilizing $entropy$, a concept from information theory, to quantify $citation$ $diversity$ in scholarly literature. The underlying idea is that an entropy measure allows us to capture the evenness of citation distribution across a set of publications (or, institutions); so higher entropy values illustrate higher diversity and lower inequality of the citation distributions, and vice versa. To be more specific, high entropy signifies more even distribution of citation, while low entropy indicates a concentrated distribution with a few highly-cited papers. In general, $diversity$ means how spread out individuals are within a group with respect to a certain trait, that can be easily categorized or measured by $numbers$. It may be mentioned here that assessing $diversity$ within a population is a crucial issue across various applied sciences, such as ecology, biology, economics, sociology, physics, and management sciences; see, e.g., [10, 11, 12, 13].
@@
In the literature, we mostly find the use of classical Shannon entropy as the entropy-based diversity measure in different contexts [14, 15, 16]. However, in this study, we use a new measure for diversity quantification, based on the concept of $logarithmic$ $norm$ $entropy$ [17, 18], to analyze the citation diversity across a range of scenarios. Firstly, we analyze the citation diversity of top-ranked institutions in different countries around the world, which allows us to explore potential geographical variations in citation diversity. Then, we extend our investigation to analyze the diversity of citations received by the publications of top award winning scientists (Nobel prize winners, Abel winners and Turing award winners). This analysis will encompass various disciplines, ensuring a holistic understanding of citation patterns in research publications across different academic fields. We also explore the time evolution of the citation diversity of the award winning scientists by analysing the diversity of recent and century old award winners across various disciplines. Finally, we dis-aggregate our findings by gender, enabling a nuanced exploration of potential gender-based disparities in citation practices of various academic disciplines. This entropy-based multifaceted approach on citation diversity offers a detailed picture, capturing the subtlety and variations in citation distribution within the scientific landscape.
@@
This research facilitating an understanding of $citation$ $diversity$ in scholarly literature is presented in a clear and logical structure. In Section II, we outline the data sources for citation information and meticulously describe the data employed in the study. We provide details regarding the selection criteria for the award winning scientists, the specific disciplines included, and the identification process for top institutes across different countries. Section III serves as the foundation of our analysis. We provide a step-by-step analytical framework in developing the concept of general class of logarithmic norm entropy and its use as a measure of citation diversity. Section IV presents the findings of our investigation where we delve into the analysis of citation diversity across various scenarios and the significance of our findings are also presented in this section. Finally a concise summary of our key findings are provided in Section V, describing the main takeaways from the analysis of citation diversity across disciplines, institutions, and publication references. Furthermore, we have discussed the broader implications of our research and potential avenues for future inquiry.
@@ II Data Description
The ‘Ranking Web of Universities’ (also commonly known as the Webometrics) [19] is a comprehensive academic ranking system established in 2004, which appears twice per year since 2006. This public resource, developed by the Cybermetrics lab, encompasses over 31,000 higher education institutions (HEIs) across more than 200 countries. Webometrics employs a mix of webometric (all missions) and bibliometric (research mission) indicators to assess university performance, promoting open access to scholarly knowledge. [19] provides the detailed citation data of the top Institutes/ universities across the world through ‘Transparent Ranking: Top Universities by Citations in Top Google Scholar profiles’ [20]. We used the January 2024 edition of this data (retrieved on April 1, 2024) for our citation analysis. The detailed data of the citation counts of top institutes can be found in [19].
Moreover, our study also leverages data from ‘Scopus’ [21], a extensive bibliographic database of peer-reviewed literature. This resource provides publication and citation information for award-winning authors across various disciplines. This unified data source allows for robust comparisons and minimizes potential biases arising from using disparate data sources. To ensure consistency, this data source is utilized for all analyses within this study. We obtained total citation data for 21 scientists in each discipline from ‘Scopus’ [21] on May 23, 2024. Additionally, we collected publication and citation data for individual scientists, including 30 Nobel laureates in physics, chemistry, and physiology/medicine (split evenly between recent and century old awardees), 15 recent Abel prize winners in mathematics, 15 recent Turing award winners in computer science, and 15 recent Nobel laureates in economics from the same source on May 23, 2024.
@@ III Methodology
III.1 General Class of logarithmic norm entropy (LNE) and diversity measure (D)
It was long known that the potential families of entropies can be used as generalised diversity measures [22]. Recently, the concept of logarithmic norm entropy (LNE) has been introduced in [17] as a new measure for quantifying diversity, justified by its better statistical efficiency and robustness properties compared to other existing classes of entropy based diversity measures. Building upon the established concept of Shannon entropy [23] and Renyi entropy [24], the LNE offers a scale-invariant generalization of the latter [18]. In this study, we leverage LNE to quantify citation diversity in scholarly literature. This unique approach allows us to assess the robustness of our findings and gain a more complete understanding of citation diversity patterns.
The generalized classes of Shannon entropy and Renyi entropy are defined, respectively, as:
$H_{\beta}^{(S)}(p)=-\frac{\sum_{i}p_{i}^{\beta}\log p_{i}}{\sum_{i}p_{i}^{\beta}},$ (1)
$H_{(\alpha,\beta)}^{(R)}(p)=\frac{1}{1-\alpha}\log{\left[\frac{(\sum_{i}p_{i}^{\alpha+\beta-1})}{(\sum_{i}p_{i}^{\beta})}\right]},$ (2)
where $p_{i}$ is the estimated normalized probability or frequency of the occurrence for the $i$-th group/class among the M possible mutually exclusive groups, for each $i=1,\dots M$, and $\alpha,\beta>0$ are two parameters (constants). Clearly, at $\beta=1$, Eq. (1) and Eq. (2) coincides with the classical Shannon entropy and Renyi entropy, respectively.
@@ III.1 General Class of logarithmic norm entropy (LNE) and diversity measure (D)
$\lim_{\alpha\rightarrow 1}H_{(\alpha,1)}^{(R)}(p)=\lim_{\alpha\rightarrow 1}H_{(\alpha,1)}^{(LN)}(p)=H_{1}^{(S)}(p)=-\sum_{i}p_{i}\log p_{i}.$
Imagine a finite set of $M$ categories representing diverse perspectives, denoted by $S={s_{1},...,s_{M}}$, such as citations across top institutes, the author’s own publications, gender, and discipline-wise variations, where the (discrete) probability distribution of citing a particular perspective in these categories is represented as $p=(p_{1},...,p_{M})$, with $p_{i}$ signifying the probability of a publication belonging to category $s_{i}$ for each $i=1,...,M$. Formally, we define the LNE based diversity measure (in percentage) for such a probability distribution, $p$, given by
@@
Additionally, it is noteworthy that the maximum value of the diversity measure equals $100\%$ for all members of the LNE family regardless of the values of the tuning parameters ($\alpha,\beta$). It always lies between 00 to $100$ (both inclusive) with higher values indicating greater diversity (and hence lower inequality) and vice versa.
The citation diversity measure (D), based on LNE, will be computed for each country and disciplinary group of prize winning scientists, as well as for individual award winners, replacing $p$ with its estimates $\widehat{p}$ derived from empirical data. We will also compute these metrics separately among males and females scientists within the award winning cohort.
@@ III.2 Asymptotic standard error and confidence interval
Since we are estimating the LNE based diversity measures based on empirical data, we must additionally quantify the possible extent of statistical errors associated with our estimates to draw more effective conclusions. As proved in [17], such estimates of the diversity measure (D) will be $\sqrt{n}$-consistent and asymptotically normal with the asymptotic variance being $\frac{\sigma_{(\alpha,\beta)}^{2}(p)}{n(\log M)^{2}}$, where
@@
with the notation $W_{c}(p)=\sum_{i=1}^{M}p_{i}^{c}$ for any $c>0$. Note that, $\sigma_{(\alpha,\beta)}^{2}(p)$ is symmetric in the choice of $(\alpha,\beta)$, as intuitively expected from similar behavior of the LNE measure itself. Since $\sigma^{2}_{H}(p)$ varies continuously in the citation distribution ($p$), we can reliably estimate it using our empirical data by replacing $p$ by its estimates $\widehat{p}$. Finally, taking square root of the estimated asymptotic variances, we get the (asymptotic) standared error (say $s$) of the estimated diversity measure (D), with lower values indicating more reliable diversity estimates.
By utilizing the standard errors (s) of the estimated diversity (D) in all our cases, we have additionally computed and plotted the $95\%$ confidence intervals for the diversity measures as given by $(D-1.96s,D+1.96s)$. This formula is obtained from the standard theory of statistical inference by utilizing the asymptotic normality of the diversity estimate. Note that, the length of the confidence interval is directly proportional to the standard error, and hence indicates the reliability of the estimated diversity; shorter the confidence interval more reliable our estimates are. Moreover, such confidence intervals also help us to statistically compare the diversity measure for two contexts (e.g., countries, subjects, or scientists); the two diversity values can be inferred to be significantly different at $5\%$ level if the associated $95\%$ confidence intervals do not overlap. This gives us a simple visual way to identify contexts having statistically similar or dissimilar diversities by just comparing the plots of their confidence intervals as presented in the following sections. Throughout our entire analysis, we have used $\alpha=2.0$ and $\beta=0.5$ in the definition of the LNE based diversity to achieve lower (asymptotic) standard error, and hence the shorter confidence interval, based on the illustrations provided in [17].
@@ IV Results and Discussions
IV.1 Citation analysis among top institutes across different countries
There is an increasing interest in world wide university rankings through various metrics [25]. The rich and exhaustive data of $university$ $rankings$ motivates one to analyse it from different angles and various perspectives [26]. In the present investigation of $citation$ $diversity$ we consider the ranking of the universities/ institutes, ($N_{I}$), according to their total number of citations ($N_{cit}$). This helps us in examining the distribution pattern of total citations, across all disciplines, of the world wide top institutions of a country as well as the citation diversity measure (D) among the top institutions of each country. This study helps in delineating the geographical variation [27] of research activities.

### v2

@@ Abstract
This study explores the citation diversity in scholarly literature, analyzing different patterns of citations observed within different countries and academic disciplines. We examine citation distributions across top institutions within certain countries and find that the higher end of the distribution follows a Power Law or Pareto Law pattern; the scaling exponent of the Pareto Law varies depending on the number of top institutions included in the analysis. By adopting a novel entropy-based diversity measure, our findings reveal that countries with both small and large economies tend to cluster similarly in terms of citation diversity. The composition of countries within each group changes as the number of top institutions considered in the analysis varies. Moreover, we analyze citation diversity among award-winning scientists across six scientific disciplines, finding significant variations. We also explore the evolution of citation diversity over the past century across multiple fields. A gender-based study in several disciplines confirms varying citation diversities among male and female scientists. Our innovative citation diversity measure stands out as a valuable tool for assessing the unevenness of citation distributions, providing deeper insights that go beyond what traditional citation counts alone can reveal. This comprehensive analysis enhances our understanding of global scientific contributions and fosters a more equitable view of academic achievements.
Keywords: Citation diversity; Diversity measure; Logarithmic norm entropy; Scholarly literature; Award winners.
@@ I Introduction
Citations are the currency of academia, reflecting impacts and influences of research publications. Ideally, a fair citation landscape would see recognition distributed proportionally to the quality and contribution of research. The term ‘quality’ generally refers to the rigor, originality, and reliability of the research, while ‘contribution of research’ refers to the significance or impact the research has within and beyond academia on advancing knowledge, solving problems, or influencing a field of study. Both factors play a crucial role in determining why certain papers receive high citation counts. Nevertheless, citations are also influenced by external factors such as visibility, collaboration networks, and research trends, rather than just intrinsic quality and contribution.
However, in practice, citations are often unevenly distributed, with a small number of papers receiving a disproportionately large share of citations, while the majority receive far fewer– a phenomenon widely discussed as citation inequality [17, 13, 37, 15]. Measuring the inequality in citation patterns has been a central focus of bibliometric studies, with many approaches borrowing from economic inequality metrics [9, 7]. There are numerous studies [10, 6, 30, 45, 39, 38, 26] on citation inequality using various inequality indices originally developed in economics (like Gini index), as well as entropy-based measures (like Shannon entropy) [33, 36]. Moreover, indices like the Hirsch index (h-index) have been utilized to summarize citation distributions [34, 21, 24, 49].
$Citation~{}diversity$, on the other hand, measures the variety and evenness of the distribution of citations across different categories, which may be institutes, authors, disciplines, etc. [46, 12, 35]. Unlike inequality, which focuses on how citations are numerically distributed, diversity captures the breadth of influence a paper or institution has across multiple fields. High citation diversity indicates a wide-ranging impact over various areas, while low diversity suggests influence being concentrated within a narrow domain. It may be mentioned here that assessing diversity within a population is a crucial issue across various applied sciences, such as ecology, biology, economics, sociology, physics, and management sciences; see, e.g., [14, 32, 29, 47, 31]. However, the potential of generalized entropy measures, which offer a versatile framework for assessing diversity, has not been fully explored in the context of citation diversity. This gap presents an opportunity to use novel entropy measures for examining the breadth and evenness of citation distribuhavetions.
@@
This study introduces a two-parameter generalization of the Renyi entropy to measure citation diversity, emphasizing the evenness of citations across different categories rather than focusing on citation counts alone. The methodology builds on the concept of logarithmic norm entropy [20, 19], adapted from information theory to quantify how evenly citations are spread within research domains, institutions, or disciplines. Higher entropy reflects a broader and more uniform citation distribution, while lower entropy indicates concentration within fewer categories. We apply this framework to explore variations in citation diversity globally, beginning with top-ranked academic institutions across different countries to explore potential geographical variations. Then, we extend our investigation to analyze the diversity of citations received by publications of top award winning scientists (Nobel prize winners, Abel winners and Turing award winners). This analysis will encompass various disciplines, ensuring a holistic understanding of citation patterns in research publications across different academic fields. We also explore the time evolution of the citation diversity of the award winning scientists by analyzing the diversity of recent and century old award winners across various disciplines. Finally, we dis-aggregate our findings by gender, enabling a nuanced exploration of potential gender-based disparities in citation practices of various academic disciplines. This entropy-based multifaceted approach on citation diversity offers a detailed picture, havecapturing the subtlety and variations in citation distribution within the scientific landscape.
@@
This research facilitating an understanding of $citation$ $diversity$ in scholarly literature is presented in a clear and logical structure. In Section II, we outline data sources for citation information and meticulously describe the data employed in the study. We provide details regarding the selection criteria for the award winning scientists, the specific disciplines included, and the identification process for top institutes across different countries. Section III serves as the foundation of our analysis, providing a step-by-step analytical framework in developing the concept of general class of logarithmic norm entropy and its use as a measure of citation diversity. Section IV presents the findings of our investigation where we delve into the analysis of citation diversity across various scenarios and the significance of our findings is also presented therein. Finally a concise summary of our key findings are provided in Section V, describing the main takeaways from our analyses. Furthermore, we have also discussed the broader implications of our research and potential avenues for future inquiry.
@@ II Data Description
The ‘Ranking Web of Universities’ (also commonly known as the Webometrics) [1] is a comprehensive academic ranking system established in 2004, which appears twice per year since 2006. This public resource, developed by the Cybermetrics lab, encompasses over 31,000 higher education institutions or universities (referred to as the HEIs) across more than 200 countries. Webometrics employs a mix of webometric (all missions) and bibliometric (research mission) indicators to assess university performance, promoting open access to scholarly knowledge. It provides the detailed citation data of the top HEIs across the world through ‘Transparent Ranking: Top Universities by Citations in Top Google Scholar profiles’ [2]. We used the January 2024 edition of this data (retrieved on April 1, 2024) for our citation analysis. The detailed data on citation counts of top institutes can be found in [1].
Moreover, our study also leverages data from ‘Scopus’ [3], an extensive bibliographic database of peer-reviewed literature. We have used this resource for obtaining publication and citation information for award-winning authors across various disciplines. This unified data source allows for robust comparisons and minimizes potential biases arising from using disparate data sources. To ensure consistency, we obtain total citation data for 21 scientists in each discipline from ‘Scopus’ [3] on May 23, 2024. Additionally, we collect publication and citation data for individual scientists, including 30 Nobel laureates in physics, chemistry, and physiology/medicine (split evenly between recent and century old awardees), 15 recent Abel prize winners in mathematics, 15 recent Turing award winners in computer science, and 15 recent Nobel laureates in economics from the same source [3] on May 23, 2024.
@@ III Methodology
III.1 The General Class of logarithmic norm entropy (LNE) and diversity measure (D)
It was long known that potential families of entropy measures can be used as generalized diversity measures [40]. Recently, the concept of logarithmic norm entropy (LNE) has been introduced in [18] as a new measure for quantifying diversity, justified by its better statistical efficiency and robustness properties compared to other existing classes of entropy based diversity measures. Building upon the established concept of Shannon entropy [44] and Renyi entropy [42], the LNE offers a scale-invariant generalization of the latter [19]. In this study, we leverage LNE to quantify citation diversity in scholarly literature. This unique approach allows us to assess the robustness of our findings and gain a more complete understanding of citation diversity patterns.
Consider a finite set of $M$ categories (denoted as $c_{1},...,c_{M}$) representing different domains of applications, such as citation patterns across top institutes, the author’s own publications, gender-based differences, and discipline-wise variations. The probability distribution over these categories is represented as $p=(p_{1},...,p_{M})$, where $p_{i}$ signifies the probability associated with category $c_{i}$ for each $i=1,...,M$. These probabilities are normalized to have the sum equal to one. The value of $M$, the number of categories, depends on the specific context of the analysis. The general classes of Shannon entropy and Renyi entropy are defined [19], respectively, as:
$H_{\beta}^{(S)}(p)=-\frac{\sum_{i}^{M}p_{i}^{\beta}\log p_{i}}{\sum_{i}p_{i}^{\beta}},~{}~{}~{}\text{where}~{}~{}\beta\in\mathbb{R}^{+},$ (1)
$H_{(\alpha,\beta)}^{(R)}(p)=\frac{1}{1-\alpha}\log{\left[\frac{(\sum_{i}p_{i}^{\alpha+\beta-1})}{(\sum_{i}p_{i}^{\beta})}\right]},~{}~{}~{}\text{where}~{}~{}\alpha,\beta\in\mathbb{R}^{+}$ (2)
Clearly, at $\beta=1$, Eq. (1) and Eq. (2) coincides with the classical Shannon entropy and the Renyi entropy, respectively.
Several other one and two-parameter generalizations of the entropy functional have been introduced in the literature, though their practical relevances and experimental validity vary. One notable example is a generalization of Renyi entropy, known as the Kapur’s generalized entropy [27, 28] of order $\alpha$ and type $\beta$, which is defined as:
$H_{(\alpha,\beta)}^{(K)}(p)=\frac{1}{\beta-\alpha}\log{\left[\frac{\sum_{i}p_{i}^{\alpha}}{\sum_{i}p_{i}^{\beta}}\right]},~{}~{}~{}\text{where}~{}~{}\alpha,\beta\in\mathbb{R}^{+}$ (3)
In this case also, when $\beta=1$, Eq. 3 coincides with the Renyi entropy. This is a two-parameter generalization of Renyi entropy but it is important to note that neither the Renyi nor the Kapur’s generalized entropy measures are scale-invariant.
@@ III.1 The General Class of logarithmic norm entropy (LNE) and diversity measure (D)
$\lim_{\alpha\rightarrow 1}H_{(\alpha,1)}^{(R)}(p)=\lim_{\alpha\rightarrow 1}H_{(\alpha,1)}^{(K)}(p)=\lim_{\alpha\rightarrow 1}H_{(\alpha,1)}^{(LN)}(p)=H_{1}^{(S)}(p)=-\sum_{i}p_{i}\log p_{i}.$
Clearly, the maximum value of the diversity measure (D) equal $100\%$ for all members of the LNE family regardless of the values of the tuning parameters ($\alpha,\beta$). It always lies between 00 to $100$ (both inclusive) with higher values indicating greater diversity and vice versa. The citation diversity measure (D), based on LNE, will be computed for each country and disciplinary group of prize winning scientists, as well as for individual award winners, replacing $p$ with its estimates $\widehat{p}$ derived from empirical data. We will also compute these metrics separately among males and females scientists within the award winning cohort.
Following (4) and noting that the maximum possible value of all these entropies is $\log M$ for a model with $M$ categories, one can define the general Diversity measure (expressed in percentage for convenience) as:
@@
For computation of $D$ in different such domains of applications, $M$ takes different values. For example, when analyzing top institutes, $c_{1},c_{2},\ldots,c_{M}$ represents the set of institutes, and $M$ is the number of institutes considered within each country (i.e., M=10 while studying citation diversity among top 10 institutes, and so on) and the citation diversity, $D$, is then computed based on the citation data of these $M$ institutes for each country. The number of countries considered is dependent on the data availability and hence it varies depending on the ranking level. For top 10 institutes, we could include 72 countries, based on data availability, while for the study of top 20 institutes we could only use data from 55 countries, and similarly 25 countries for analyzing diversity among top 50 institutes.
@@ III.2 Asymptotic standard error and confidence interval
Since we are estimating the LNE based diversity measures from empirical data, we must additionally quantify the extent of statistical errors associated with our estimates to draw more effective conclusions. As proved in [20], such estimates of the diversity measure (D) will be $\sqrt{n}$-consistent and asymptotically normal with the asymptotic variance being $\frac{\sigma_{(\alpha,\beta)}^{2}(p)}{n(\log M)^{2}}$, where
@@
with the notation $W_{c}(p)=\sum_{i=1}^{M}p_{i}^{c}$ for any $c>0$. Note that, $\sigma_{(\alpha,\beta)}^{2}(p)$ is symmetric in the choice of $(\alpha,\beta)$, as intuitively expected from similar behavior of the LNE measure itself. Since $\sigma^{2}_{(\alpha,\beta)}(p)$ varies continuously in the citation distribution ($p$), we can reliably estimate it using our empirical data by replacing $p$ by its estimates $\widehat{p}$. Finally, taking square root of the estimated asymptotic variances, we get the (asymptotic) standard error (say $s$) of the estimated diversity measure (D), with lower values indicating more reliable diversity estimates.
By utilizing the standard errors (s) of the estimated diversity (D) in all our cases, we have computed and plotted the $95\%$ confidence intervals for the diversity measures as given by $(D-1.96s,D+1.96s)$. This formula is obtained from the standard theory of statistical inference by utilizing the asymptotic normality of the diversity estimate. Note that, the length of the confidence interval is directly proportional to the standard error, and hence indicates the reliability of the estimated diversity; shorter the confidence interval more reliable our estimates are. Moreover, such confidence intervals also help us to statistically compare the diversity measures for two contexts (e.g., countries, subjects, or scientists); two diversity values can be inferred to be significantly different at $5\%$ level if the associated $95\%$ confidence intervals do not overlap. This gives us a simple visual way to identify contexts having statistically similar or dissimilar diversities by just comparing the plots of their confidence intervals as presented in the following sections. Throughout our entire analysis, we have used $\alpha=2.0$ and $\beta=0.5$ in the definition of the LNE based diversity. Although there exists a whole class of LNE measures with different choices of $\alpha$ and $\beta$ tailored to various applications, this specific combination is found to provide the most meaningful results for our citation data, along with having lower standard errors and narrower confidence intervals; it was also recommended in [20] from statistical considerations.
@@ IV Results and Discussions
IV.1 Citation analysis among top institutes within different countries
There is an increasing interest in global university rankings through various metrics [5]. The rich and exhaustive data of $university$ $rankings$ motivates one to analyse it from different angles and various perspectives [11]. In the present investigation of $citation$ $diversity$ we consider the ranking of the universities/ institutes, ($N_{I}$), according to their total number of citations ($N_{cit}$). This helps us in examining the distribution pattern of total citations, across all disciplines, of top institutions of a country as well as the citation diversity measure (D) among the top institutions of each country. This study helps in delineating the geographical variation [22] of research activities.